OpenStudy (anonymous):

. Write an equation in point-slope form of a line parallel to the line y = 2x - 6 and passes through the point (2, 3).

Directrix (directrix):

Parallel lines have the same slope. y = 2x - 6 is already written in slope-intercept form. The slope is 2. Therefore, the slope of the line passing through the point (2,3) will be 2. Using the point-slope form of writing the equation of a line, y - y1 = m (x - x1) y - 3 = 2 (x - 2) y - 3 = 2x - 4 y = 2x -4 + 3 y = 2x - 1 ---> Answer